OpenStudy (anonymous):

I have this question and I don't quite know how to solve the equation, because there is both y and x. So: x + 3 / 3 = y + 2 / 2

OpenStudy (anonymous):

and then it says, "if x + 3 / 3 = y + 2/2, then x / 3 = _____"

OpenStudy (anonymous):

And here is what I did: 2 ( x + 3), which gave me 2x + 6 = 3 (y + 2), which gave me 2x + 6 = 3y + 6, now I don't know what to do.

OpenStudy (anonymous):

so is (x +3) and (y + 2) in parentheses in the original equation?

OpenStudy (anonymous):

No, I was just distributing :/

OpenStudy (anonymous):

It looks like this

OpenStudy (phi):

I would write it as \[ \frac{x+3}{3}= \frac{x}{3}+ \frac{3}{3}= \frac{x}{3}+1 \]

OpenStudy (phi):

and do the same thing to the other side with y

OpenStudy (anonymous):

I don't even understand why it is says then x/3 = ____. Like what is it asking me to do ?

OpenStudy (phi):

(y+2)/2 = y/2 + 1 you have \[ \frac{x}{3}+1= \frac{y}{2} + 1 \] subtract 1 from both sides \[ \frac{x}{3}= \frac{y}{2} \]

OpenStudy (phi):

*** I don't even understand why it is says then x/3 = ____. Like what is it asking me to do ? *** they are asking you to "get x/3 by itself" on one side of the equation once you do that, the answer is whatever is on the other side.

OpenStudy (anonymous):

okay so let me go back, alright, yes I see what you are doing.

OpenStudy (anonymous):

X + 3 / 3 = x /2 + 3/3 = x / 3 + 1 (y + 2) / 2 = y + 2 + 1 I subtracted the 1 and I got = x / 3 = y/ 2 Is this right? I know it looks lik eyours, but I actually worked it out, and I don't think I would've gotten it ever ina million years without guidance. I am not very good when I don't understand what is being asked (second language and all)

OpenStudy (phi):

I assume the 2 in X + 3 / 3 = x /2 + 3/3 is a typo and you mean x/3 ?? (y+2)/2 = y+2+1 is way off. typo or you don't know how to do that?

OpenStudy (anonymous):

Major typo. I got it correct though, I just spoke with my teacher. Sorry about the typo but I got it now :) Didn't see it.